Please please help me. I tried on my own but I don't know if I did it right and i don't want to look stupid. My professor has already been more than patient with me.
solve the following equations:
#1
A.
square root x-2=1
(show work)
****heres what I have*****
(sqrt x-2)^2=(1)^2
x-2=1
x=3
B.
sqrtx^3=27
show work
********heres what I have******
(sqrtx^3)^2=(27)^2
x^3=729
C.
3sqrtx^2=9
I really tried here but couldn't figure it out.
I appreciate your time.

Answer by checkley71(8403)

SQRT(X-2)=1 OR SQUARING BOTH SIDES WE GET X-2=1 OR X=3
PROOF SQRT3-2=1 OR SQRT1=1 1=1
SQRT(X~3)=27 OR SQUARING BOTH SIDES WE GET X~3=27*27 OR X~3=3*3*3*3*3*3 OR
X*X*X=9*9*9 THUS X=9